OpenStudy (kaliyah0524):

Medal! (x-b)(x+b)(x^2-b^2)

OpenStudy (danjs):

Do the first two first.. \[\large (x-b)(x+b) = x^2 +bx - bx - b^2 = x^2 - b^2\] so it becomes \[\large (x^2 - b^2)(x^2 - b^2)\]

OpenStudy (kaliyah0524):

So would the answer be x^4 - b^4x^4 +b^4 @DanJS

OpenStudy (danjs):

\[(x^2-b^2)(x^2-b^2)=x^2*x^2-x^2*b^2-b^2*x^2+b^2*b^2\]

OpenStudy (danjs):

\[\large x^4 - 2x^2b^2+b^4\]

OpenStudy (kaliyah0524):

Thanks